Please help me solve this-
1/2 devided (2/3 devided 3/4) - (4/5 * 5/6)
I can't figure it out. I have the answer but I can't get the answers to match.

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

(1/2) / ((2/3) / (3/4)) - (4/5 * 5/6)
= (1/2)* [(3/4)/(2/3)] - (4/6)
= (1/2)* (9/8) - (2/3)
= (9/16)-(2/3)
= (27-32)/48
= --5/48
============
Cheers,
Stan H.

Answer by checkley71(8403)

(1/2)/([2/3]/3/4)-(4/5*5/6)
(1/2)/(2/3)*(4/3)-20/30
(1/2)/(8/9)-2/3
1/2*9/8-2/3
9/16-2/3
(9*3-2*16)/48
(27-32)/48
-5/48 answer.